If the perimiter of a rectange is 30m and the length is 3m more than twice the width what are the dimensions? Length and width. I think the length is 12 and the width is 3
Answer by mananth(16949)

If the perimiter of a rectange is 30m and the length is 3m more than twice the width what are the dimensions? Length and width. I think the length is 12 and the width is 3
let the width be x
length will be 2(x+3)
Perimeter = 2L + 2W
2x+ 2*2(x+3)= 30
2x+4x+12=30
6x+12=30
6x= 18
x= 3 wich is the width
Length = 2(x+3)
Length will be 12
You are right